OpenStudy (coledrf):

1. -(6)^-1 (1 point)

OpenStudy (anonymous):

\[a ^{b}=\frac{ 1 }{ a ^{-b} }\]

OpenStudy (anonymous):

That simplifies to..\[-\frac{ 1 }{ 6^1 }\]..or\[-\frac{ 1 }{ 6 }\]

OpenStudy (coledrf):

yay thank you i have another if you are up for it

OpenStudy (anonymous):

sure

OpenStudy (coledrf):

simplify (5)^-5(5)^7

OpenStudy (coledrf):

u still there?

OpenStudy (anonymous):

\[5^{-5(5)^7}=\frac{ 1 }{ 5^{390625} }\]

OpenStudy (anonymous):

yeah, i felt like i was messing that one up

OpenStudy (coledrf):

ty are you up for another?

OpenStudy (anonymous):

why not :P